if 2 meters of fabric costs #3.45, what should 7 meters cost?
a 16 ounce box of laundry detergent costs $2.49, how many ounces should be in a box marked $1.15?

Answer by rfer(16322)

2/3.45=7/x
2x=24.15
x=12.08
---------------
16/2.49=x/1.15
2.49x=18.4
x=7.39